# Parameter-Efficient Continual Fine-Tuning: A Survey

### v1

@@ 5.2 Performance Analysis and Trade-offs
The problem of correctly evaluating the performance of CL algorithms has been present since the early developments of the field. It is clear that accuracy alone is not enough: most of the times, standard “offline” learning models have an higher accuracy than CL solutions, but this drop may be acceptable if we measure other, but nonetheless equally important benefits brought by CL algorithms.
This is particularly evident in the context of PECFT methods. Indeed, we mentioned multiple times throughout this work that the principal concern about PTMs lies in their training and inference inefficiency, and in the large resources needed for fine-tuning. This means that, when dealing with very large architectures, pure accuracy cannot be the only metrics employed to evaluate such models.
Firstly, in dynamic scenarios the accuracy on the last task only is not so informative, since we are interested in retaining knowledge and performance on all seen task. Formally, we define the Average Accuracy [12] as:
@@
Apart from pure accuracy, the CL fields principally studies approaches to diminish the catastrophic forgetting phenomenon, hence it comes straightforwardly that it represents an important measure to keep into consideration.
@@
where $\text{where }f_{k,j}=\max$ (${}_{l\in\{1,\dots,k-1\}}(a_{l,j}-a_{k,j}),\quad\forall j<k$
Given the sequential nature of the learning scenarios, it is useful to measure what is the influence of learning multiple tasks, i.e. how much training on a task affects the performance on other tasks. This is particularly interesting because it gives practical insights on how fast the model learns through time. For this reasons, forward transfer and backward transfer [90, 26] are introduced.
@@ 6.1 Multi-modality Pre-trained Models
Different pre-trained models’ influence on CL methods’ effectiveness requires deeper analysis. Research not only indicates variations in performance between models trained from scratch and those utilizing pre-trained architectures but also shows that different pre-trained models benefit differently from existing continual learning approaches. [72] Thus, exploring the selection of the optimal pre-trained model and its best-fitted continual learning strategy given a real-world task can lead to meaningful contributions to guiding industry practice. Additionally, while a majority of the experiments were tested over visual and natural language modalities, video and audio are domains where the application of continual learning can be valuable to explore as well. [102, 92]
@@ 6.2 Model Merging
Model merging presents an exciting opportunity in CL, by combining multiple expert models, each specialized in different aspects of a task, we can create a system that not only mitigates issues like catastrophic forgetting but also benefits from the diverse strengths of each model. This becomes particularly important in dynamic, evolving domains, as it allows the model to expand its knowledge over time without forgetting what it has previously learned. However, a common problem that these model merging solutions, such as Task Arithmetic [55], often encounter is parameter interference, which leads to significant performance degradation when these expert models are merged. Some works such as TIES-MERGING [148] and DARE [152] have led to significant improvements in model merging. [148] addresses interference by resetting parameters that have only changed minimally, resolving sign conflicts, and merging only those parameters that align with the final agreed-upon sign. [152], on the other hand, eliminates redundant delta parameters by randomly dropping them and rescaling the remaining ones, which has shown tremendous effectiveness in sparsifying and merging multiple expert models without significant performance loss.
Typical model merging scenarios often require combining pre-existing expert models—each specialized in a specific task—into one unified system. However, this static approach falls short in scenarios where new tasks emerge over time. In continual learning, we face the challenge of incrementally integrating new task-specific models without retraining the entire system. Recent advances in dynamic model merging address this by tackling issues such as parameter interference, memory efficiency, and sequential integration, enabling systems that adapt more effectively as new tasks are encountered. For instance, MagMax[99] introduces a framework that merges task-specific models using sequential fine-tuning combined with a maximum magnitude weight selection strategy. This approach integrates new information effectively and preserves the integrity of earlier learning to help tackle CT. In contrast, Representation Surgery for Multitask Model Learning [149] addresses a different challenge. Here, the focus is on mitigating the representation bias that emerges when merging models trained on disparate tasks. By inserting a lightweight, task-specific module—dubbed “Surgery”—the method realigns the merged model’s internal representations with those of the individual models, thereby enhancing overall performance in multitask scenarios. While Adaptive LoRA Merging for Domain Incremental Learning[20] highlights the limitations of fixed-weight merging by proposing an adaptive mechanism that dynamically computes merging coefficients. This flexibility allows the system to balance the contributions of new and old domains, ensuring robust performance in evolving environments while reducing manual tuning.
@@ 6.3 From Classification to Reasoning
Large models have been considered to exhibit “emergent” phenomena, yet they still have ways to go before they can effectively handle more complex reasoning tasks. [145, 121] CL is crucial in this regard and will also face new challenges.
The evolution of tasks from single-ability tasks, such as image classification, to more complex reasoning tasks demands a re-evaluation of existing continual learning methods or the proposal of new approaches. Classically, many methods in CL are evaluated using classification datasets such as CIFAR and ImageNet. There has been a recent push towards transitioning from hard split datasets to benchmarks that reflect more natural temporal and task shifts. [81, 135] Despite this, very few studies address the specific requirement for more complex reasoning tasks, such as visual question answering [100, 150, 74], decision-making [27], motion programming [70], scene understanding [89].
Reasoning tasks can differ significantly in their requirements. For instance, in image classification, a model continually improves its ability to recognize objects within a single modality and primarily focuses on learning representations for each class. However, in reasoning tasks like visual question answering, a model in many cases need to acquire multiple new abilities such as attribute recognition, knowledge reasoning based on changing demands. [74, 158]
Representation methods that focus on retain representative embedding from historical task typically fail to handle this situation, thus cannot be applied directly. Traditional replay techniques may prove insufficient in addressing these more intricate tasks as well. Some good examples that dive deeper into the specific need for reasoning tasks are: [98] maps sub-symbolic inputs to high-level concepts and explicitly replay concepts for CL in neuro-symbolic(NeSy) architectures such as DeepProbLog [97];  [74] utilizes scene graph as a prompt to represent previous images in visual question answering task and replays pseudo scene graphs alongside corresponding QA pairs.
@@ 6.4 More Realistic and Practical Settings
Future research should explore Continual Learning in more practical and realistic settings. This includes scenarios characterized by limited computational budgets. A recent survey [134] pointed out that existing continual learning methods tend to consider memory constraint, but computational costs are not extensively considered. While the cost of memory and privacy concerns may seem prohibitive, in practice, storing large datasets for extended periods can relatively inexpensive compared to the computational costs of training models on such datasets. [107] This is especially relevant for methods relying on replay mechanisms.

### v2

@@ 5.5 Performance Analysis and Trade-offs
The problem of correctly evaluating the performance of CL algorithms has been present since the early developments of the field. It is clear that accuracy alone is insufficient to comprehensively assess these methods in CL scenarios: traditional “offline” learning models typically achieve higher accuracy than CL solutions. However, dynamic environments introduce additional performance considerations, making it necessary to balance trade-offs among the multiple evaluation metrics. This is particularly evident in the context of PECFT methods. Indeed, we have mentioned multiple times throughout this work that the principal concern about PTMs lies in their training and inference inefficiency, as well as the large resources required for fine-tuning. This means that, when dealing with very large architectures, pure accuracy cannot be the only metric employed to evaluate such models.
Firstly, CL scenarios usually involve multiple learning tasks, each with its own training and a test sets. At first glance, the accuracy on the final task might seem like a meaningful indicator of overall model capabilities, since it represents the end of the training phase. However, a core principles of CL is the ability to retain knowledge across the entire learning experience. As such, more comprehensive accuracy metrics are required to capture this aspect effectively. Formally, we define the Average Accuracy [12] as:
@@
Apart from pure accuracy, the CL fields principally study approaches to diminish the catastrophic forgetting phenomenon; hence it comes straightforwardly that it represents an important measure to keep into consideration.
@@
$\text{where }f_{k,j}=\max$ ${}_{l\in\{1,\dots,k-1\}}(a_{l,j}-a_{k,j}),\quad\forall j<k$
Given the sequential nature of the learning scenarios, it is useful to measure what is the influence of learning multiple tasks, i.e. how much training on a task affects the performance on other tasks. This is particularly interesting because it gives practical insights on how fast the model learns through time. For this reason, forward transfer and backward transfer [96, 27] are introduced.
@@
5.6 Comparative Performance Evaluation
The current AI landscape is filled with a vast number of benchmarks and evaluation metrics, and individual methods often rely on different experimental settings. On top of that, a wide set of variables, such as the number of training epochs, the learning rate and PEFT-specific parameters, e.g. LoRA rank or prompt size, can significantly influence performance. This variability makes fair comparisons among different models very challenging, hindering the ability to determine which method performs best in which scenario and, more importantly, why. To address this, the following section presents an experimental analysis of the most representative PECFT methods, aiming to provide guidance for researchers in a performance-oriented perspective.
5.6.1 Accuracy Performance Analysis
The comparative analysis of PECFT methods reveals a complex landscape where no single approach achieves universal superiority, with performance being highly dependent on both dataset characteristics and task sequence length. While this evaluation focuses on the most widely adopted PECFT approaches rather than exhaustively covering all variants, the selected methods represent the core paradigms that have gained significant traction in the field, providing meaningful insights into fundamental trade-offs. The results of our evaluation of the different methods is available in Table 4.
Method CIFAR-100 CUB-200 Tiny Imagenet
10 Tasks 20 Tasks 10 Tasks 20 Tasks 10 Tasks 20 Tasks
L2P 80.01 ± 0.93 72.48 ± 0.33 64.32 ± 0.18 56.62 ± 0.77 80.24 ± 0.21 75.44 ± 0.87
DualPrompt 80.45 ± 0.33 77.83 ± 0.40 69.29 ± 2.73 57.77 ± 3.58 83.32 ± 0.27 78.77 ± 0.01
CODA-Prompt 83.65 ± 1.34 74.18 ± 1.03 73.23 ± 0.21 60.69 ± 1.97 89.33 ± 0.40 83.88 ± 0.27
InfLoRA 84.64 ± 0.41 78.14 ± 0.41 73.42 ± 1.06 63.61 ± 0.37 83.29 ± 0.18 80.21 ± 0.55
SD-LoRA 86.02 ± 0.24 82.08 ± 1.92 72.07 ± 3.77 68.36 ± 3.06 88.76 ± 0.65 85.32 ± 0.82
LoRAC-IPC 87.79 ± 0.19 84.59 ± 0.51 81.24 ± 0.57 72.44 ± 1.91 84.84 ± 0.27 83.43 ± 0.23
SEMA 86.41 ± 0.55 81.88 ± 0.92 74.19 ± 0.96 65.15 ± 0.95 89.07 ± 0.69 87.27 ± 0.38
APER (adapter) 84.16 ± 0.14 82.06 ± 0.23 86.31 ± 0.01 86.12 ± 0.04 76.84 ± 0.54 70.31 ± 0.20
EASE 84.29 ± 0.54 82.51 ± 0.12 81.60 ± 0.28 81.29 ± 0.07 88.39 ± 0.08 85.85 ± 0.13
LAE 84.84 ± 0.13 83.52 ± 0.12 79.77 ± 0.45 73.74 ± 0.43 85.58 ± 0.15 83.52 ± 0.12
Table 4: Average Accuracy (%, $\uparrow$) of PECFT methods after seeing all tasks. Each method is assessed across multiple datasets and learning scenarios. The best performance is highlighted in bold, while the second best result is underlined.
Adapter-based methods
Adapter-based methods demonstrate consistent performance across the different benchmarks, with APER achieving exceptional results on fine-grained classification task (86.31% on CUB-200 with 10 tasks), showing also remarkable scalability when increasing the number of tasks (-0.19% accuracy). EASE shows competitive performance, obtaining the second best results in multiple scenarios, while demonstrating a good scalability in longer scenarios. SEMA has strong performance on broader classification tasks (86.41% on CIFAR-100, 89.07% on Tiny-ImageNet), obtaining instead slightly worse results on CUB-200. It also suffers severe scalability issues, especially on CUB-200, where it has a 9.04% drop in performance when going from 10 to 20 tasks scenarios. On average, adapter-based models obtain the highest accuracy, which is expected since they typically require the higher number of trainable parameters, as seen in Table 3.
LoRA-based methods
LoRA approaches perform competitively, with LoRAC-IPC achieving the highest scores on CIFAR-100. Together with SD-LoRA, it shows consistent cross-dataset performance, with moderate but predictable scalability–3.4-3.9% degradation on longer sequences. On the other hand, InfLoRA’s subspace constraints show concerning scalability issues on complex datasets, despite superior forgetting metrics. Overall, LoRA-based methods are reliable choices for diverse scenarios, also considering their increased efficiency compared to adapter-based architectures.
Prompt-based methods
Traditional prompt-based methods face fundamental limitations, primarly due to the limited learning capacity resulting from the small number of additional trainable parameters they involve. Even more involved variants like CODA-Prompt exhibit poor scalability–showing accuracy drops ranging from 9.47% to 12.54% as the number of tasks increases, and consistently lower absolute performance, lagging 3-6% behind other PECFT alternatives. In contrast, simpler methods like DualPrompt surprisingly demonstrate greater stability across different domains. In summary, prompt-based methods represent the lightest and most efficient PECFT alternatives, showing nonetheless inferior performance when compared to other approaches.
Unified frameworks
Unified frameworks like LAE show promising scalability, achieving only a 1.32% drop on CIFAR-100, indicating that these meta frameworks can effectively leverage the strengths of their underlying PEFT components. However, their evaluation within an adapter-based setting limits our understanding on how well they generalize across different PEFT strategies.
These results suggest that practitioners should prioritize adapter methods for peak performance on specific domains, LoRA methods for consistent multi-domain deployment, and consider unified frameworks as promising approaches for combining multiple PEFT strengths, while carefully weighing the scalability-performance trade-offs inherent in each approach, as the 10-to-20 task scaling reveals that method stability often matters more than initial performance in long-term continual learning scenarios.
Method Average Forgetting
L2P 10.61 ± 0.37
DualPrompt 6.43 ± 0.14
CODA-Prompt 4.98 ± 0.52
InfLoRA 4.61 ± 0.44
SD-LoRA 6.19 ± 1.09
LoRAC-IPC 5.83 ± 0.38
SEMA 6.40 ± 0.87
APER adapter 6.93 ± 0.14
EASE 6.80 ± 0.11
LAE 6.18 ± 0.19
Table 5: Average Forgetting(%, $\downarrow$) of different PECFT methods, assessed on CIFAR100 with 10 tasks. The best performance is highlighted in bold, while the second best result is underlined.
5.6.2 Forgetting Performance Analysis
The catastrophic forgetting analysis, shown in Table 5, reveals complex patterns that are surprisingly decoupled from performance metrics, challenging assumptions about method effectiveness.
Within LoRA-based approaches, InfLoRA achieves exceptional forgetting mitigation (4.61%) through subspace-constrained learning, demonstrating that orthogonal parameter updates effectively preserve previous knowledge without memory overhead. LoRAC-IPC has moderate forgetting, showing the effectiveness of their two-stage approach to fight parameter interference. SD-LoRA follows closely, indicating that decoupled approaches help but cannot match subspace constraints.
Adapter-based methods exhibit remarkably consistent moderate forgetting, with SEMA, EASE and APER clustered tightly around 6-7% forgetting, despite their varying performance levels. This suggests that adapter architectures provide inherent baseline stability but do not excel in knowledge preservation.
Lastly, prompt-based methods exhibit significant variability. CODA-Prompt achieves surprisingly strong performance (4.98%), rivaling to the best LoRA-based approaches, while L2P suffers from severe forgetting (10.61%), nearly doubling the forgetting of any other competing method. In contrast, DualPrompt demonstrates moderate stability, highlighting the critical role of the prompt design in determining method robustness.
The notable decoupling between performance and forgetting creates complex trade-offs: InfLoRA combines exceptional forgetting resistance with competitive performance, making it ideal for stability-critical applications, while high-performing methods like SEMA achieve peak accuracy at the cost of moderate forgetting, and some methods like L2P fail catastrophically on both metrics.
This analysis suggests that practitioners must explicitly prioritize between performance and stability, as no method achieves excellence in both dimensions, with subspace-constrained approaches like InfLoRA offering the best stability-performance balance for these long-term continual learning scenarios.
5.7 Core Design Choices of PECFT
PECFT often incorporate different design/architectural choices which offer several advantages and trade-offs. It is important to understand the impact of the choices that can influence the CL performance when using PEFTs. Rather than viewing PECFT methods as isolated techniques for CL, we identify some core design principles based on both empirical evidence from our comparative analysis and theoretical insights from the literature.
5.7.1 Parameter Separation
When working with PECFT, a key design choice is deciding which parameters should remain frozen and which ones should be allowed to change for new tasks. This decision is crucial to how the methods learn new tasks and also maintain old knowledge. One such example is freezing the backbone. Most successful PECFT methods preserve the pre-trained model parameters while introducing task-specific adaptations through the addition of extra components. This works mostly because pre-trained models encode general-purpose representations learned from vast datasets that capture fundamental patterns useful across multiple domains. When these representations are modified during CL, the model risks losing this general knowledge, leading to catastrophic forgetting. By keeping the backbone frozen, PECFT methods ensure that this foundational knowledge remains intact while allowing task-specific modifications to learn new knowledge through dedicated parameters.
Building on the freezing of the backbone, the next critical decision is how to introduce task-specific adaptations. While all PECFT methods preserve the pre-trained backbone, they differ fundamentally in how they add learning capacity for new tasks.
  • 1.
    Modular Separation: Introduces dedicated task-specific modules while keeping backbone frozen. This provides strong modularity but requires careful routing mechanisms.
  • 2.
    Subspace Separation: Constrains updates to low-dimensional subspaces of the original parameter space. This maintains parameter efficiency but can create interference between tasks.
  • 3.
    Input Space Separation: Modifies the input representation while keeping all model parameters frozen.
To illustrate the impact of this design choice, Figure 3 provides a visual representation of the architectures that implement different parameter separation strategies.
5.7.2 Interference Prevention
In the context of PECFT, the phenomenon of catastrophic forgetting may be caused also by interference among the different tasks. Therefore, avoiding this kind of issue requires explicit architectural design choices, which we briefly summarize in the following.
Orthogonality Constraints
One approach enforces mathematical orthogonality between task-specific updates through gradient projection techniques that constrain new learning to subspaces orthogonal to previous tasks [116]. Recent theoretical work has identified task-level feature orthogonality as a key factor influencing PECFT performance, alongside training sample size and regularization [87]. This provides the strongest theoretical guarantees against interference by ensuring that updates for new tasks cannot directly modify the directions used by previous tasks. Methods like [82, 14] enforce orthogonality between task-specific updates. This provides the strongest theoretical guarantees against interference, explaining why [82] achieves exceptional stability (4.61 ± 0.44% forgetting). However, orthogonality constraints limit the available parameter space for each task, creating potential scalability concerns as the number of tasks grows.
Figure 3: Parameter Separation Strategies
Architectural Modularity
This aims to prevent interference through physical separation of parameters, where each task uses dedicated modules while the backbone remains frozen. This architectural separation provides interference resistance through construction rather than restraint, since tasks often use entirely different parameters and they cannot interfere with each other. This includes approaches that use task specific adaptation modules with different kind of learned routing mechanism to select the adapted modules. Methods like [146, 171] show competitive results, however with performance that degrade when the number of tasks increase. In fact, while forgetting may seem reasonable on smaller tasks, this kind of separation forces modules to become very task specific. In this scenario, more tasks compete for selection, adding complexity to the selection/routing mechanism. Moreover, a wrong selection could lead to catastrophic failure especially when the task is highly challenging. For example, [146] achieve good performance (86.41% on CIFAR-100) but shows severe scalability issues on fine-grained tasks, especially when increasing the number of tasks from 10 to 20 (CUB-200: 74.19% → 65.15%, 9.04% drop).
Compositional Isolation
A third approach uses compositional mechanisms to isolate task-specific information within shared parameter spaces. This strategy defines reusable components, such as prompt vectors, that can be combined differently for different tasks through selection schemes or weighted coefficient mechanisms, i.e. creating a shared pool of reusable components that can be mixed and matched for different tasks. Instead of learning tasks with completely separate PEFT modules, or forcing the task to use the same modules with some kind of constraints, it creates a set of reusable modules where each task uses a different combination of these modules. Well-designed compositional approaches can achieve stability comparable to orthogonal methods (CODA-Prompt: 4.98% forgetting) through careful component design that minimizes conflicts between different composition strategies. Compositional methods face fundamental representational limits as task diversity increases. In fact, even sophisticated composition strategies eventually encounter situations where existing components cannot adequately represent new task requirements, leading to universal scalability degradation.
@@ 6.1 Multi-Modality Pre-trained Models
Different pre-trained models’ influence on CL methods’ effectiveness requires deeper analysis. Research not only indicates variations in performance between models trained from scratch and those utilizing pre-trained architectures but also shows that different pre-trained models benefit differently from existing continual learning approaches [74]. Thus, exploring the selection of the optimal pre-trained model and its best-fitted continual learning strategy given a real-world task can lead to meaningful contributions to guiding industry practice. Additionally, while a majority of the experiments were tested over visual and natural language modalities, video and audio are domains where the application of continual learning can be valuable to explore as well [109, 99].
@@ 6.2 Model Merging
Model merging presents an exciting opportunity in CL, by combining multiple expert models, each specialized in different aspects of a task, we can create a system that not only mitigates issues like catastrophic forgetting but also benefits from the diverse strengths of each model. This becomes particularly important in dynamic, evolving domains, as it allows the model to expand its knowledge over time without forgetting what it has previously learned. However, a common problem that these model merging solutions, such as Task Arithmetic [58], often encounter is parameter interference, which leads to significant performance degradation when these expert models are merged. Some works such as TIES-MERGING [159] and DARE [163] have led to significant improvements in model merging. TIES addresses interference by resetting parameters that have only changed minimally, resolving sign conflicts, and merging only those parameters that align with the final agreed-upon sign. DARE, on the other hand, eliminates redundant delta parameters by randomly dropping them and rescaling the remaining ones, which has shown tremendous effectiveness in sparsifying and merging multiple expert models without significant performance loss.
Typical model merging scenarios often require combining pre-existing expert models—each specialized in a specific task—into one unified system. However, this static approach falls short in scenarios where new tasks emerge over time. In continual learning, we face the challenge of incrementally integrating new task-specific models without retraining the entire system. Recent advances in dynamic model merging address this by tackling issues such as parameter interference, memory efficiency, and sequential integration, enabling systems that adapt more effectively as new tasks are encountered. For instance, MagMax [106] introduces a framework that merges task-specific models using sequential fine-tuning combined with a maximum magnitude weight selection strategy. This approach integrates new information effectively and preserves the integrity of earlier learning to help tackle catastrophic forgetting. In contrast, Representation Surgery for Multitask Model Learning [160] addresses a different challenge. Here, the focus is on mitigating the representation bias that emerges when merging models trained on disparate tasks. By inserting a lightweight, task-specific module—dubbed “Surgery”—the method realigns the merged model’s internal representations with those of the individual models, thereby enhancing overall performance in multitask scenarios. While Adaptive LoRA Merging for Domain Incremental Learning [20] highlights the limitations of fixed-weight merging by proposing an adaptive mechanism that dynamically computes merging coefficients. This flexibility allows the system to balance the contributions of new and old domains, ensuring robust performance in evolving environments while reducing manual tuning.
@@ 6.3 From Classification to Reasoning
Large models have been considered to exhibit “emergent” phenomena, yet they still have ways to go before they can effectively handle more complex reasoning tasks [154, 129]. CL is crucial in this regard and will also face new challenges.
The evolution of tasks from single-ability tasks, such as image classification, to more complex reasoning tasks demands a re-evaluation of existing continual learning methods or the proposal of new approaches. Classically, many methods in CL are evaluated using classification datasets such as CIFAR and ImageNet. There has been a recent push towards transitioning from hard split datasets to benchmarks that reflect more natural temporal and task shifts [84, 144]. Despite this, very few studies address the specific requirement for more complex reasoning tasks, such as visual question answering [107, 161, 76], decision-making [28], motion programming [72], scene understanding [95].
Reasoning tasks can differ significantly in their requirements. For instance, in image classification, a model continually improves its ability to recognize objects within a single modality and primarily focuses on learning representations for each class. However, in reasoning tasks like visual question answering, a model in many cases need to acquire multiple new abilities such as attribute recognition, knowledge reasoning based on changing demands [76, 169].
Representation methods that focus on retain representative embedding from historical task typically fail to handle this situation, thus cannot be applied directly. Traditional replay techniques may prove insufficient in addressing these more intricate tasks as well. Some good examples that dive deeper into the specific need for reasoning tasks are: [105] maps sub-symbolic inputs to high-level concepts and explicitly replay concepts for CL in neuro-symbolic architectures such as DeepProbLog [104];  [76] utilizes scene graph as a prompt to represent previous images in visual question answering task and replays pseudo scene graphs alongside corresponding QA pairs.
@@ 6.4 More Realistic and Practical Settings
Future research should explore Continual Learning in more practical and realistic settings. This includes scenarios characterized by limited computational budgets. A recent survey [143] pointed out that existing continual learning methods tend to consider memory constraint, but computational costs are not extensively considered. While the cost of memory and privacy concerns may seem prohibitive, in practice, storing large datasets for extended periods can relatively inexpensive compared to the computational costs of training models on such datasets [114]. This is especially relevant for methods relying on replay mechanisms.